### v1

# Memory-Centric Embodied Question Answer

Abstract
Embodied Question Answering (EQA) requires agents to autonomously explore and understand the environment to answer context-dependent questions. Existing frameworks typically center around the planner, which guides the stopping module, memory module, and answering module for reasoning. In this paper, we propose a memory-centric EQA framework named MemoryEQA. Unlike planner-centric EQA models where the memory module cannot fully interact with other modules, MemoryEQA flexible feeds memory information into all modules, thereby enhancing efficiency and accuracy in handling complex tasks, such as those involving multiple targets across different regions. Specifically, we establish a multi-modal hierarchical memory mechanism, which is divided into global memory that stores language-enhanced scene maps, and local memory that retains historical observations and state information. When performing EQA tasks, the multi-modal large language model is leveraged to convert memory information into the required input formats for injection into different modules. To evaluate EQA models’ memory capabilities, we constructed the MT-HM3D dataset based on HM3D, comprising 1,587 question-answer pairs involving multiple targets across various regions, which requires agents to maintain memory of exploration-acquired target information. Experimental results on HM-EQA, MT-HM3D, and OpenEQA demonstrate the effectiveness of our framework, where a 19.8% performance gain on MT-HM3D compared to baseline model further underscores memory capability’s pivotal role in resolving complex tasks.
Figure 1: This task requires the agent to identify multiple relevant entities involved in the question. Planner-centric method fail to recall that the living room sofa has already been found, and thus it searches for the sofa again, resulting in redundant exploration. In contrast, reasoning framework guided by memory chose an unexplored direction.
1 Introduction
Embodied Question Answering (EQA) [11, 14, 34, 16] is a challenging task in robotics, requiring agents to resolve questions via interaction within a dynamic and evolving environment. Existing methods [27, 39, 12, 7, 10] typically center around a planner (as shown in the left of Figure 2), coordinating modular components including memory module, answering module, and stopping module. The planner relies only on current observations to invoke the other modules. However, these methods often suffer from inefficient exploration and inaccurate response when faced with complex tasks (e.g., the tasks requires exploring multiple regions to gather information about numerous targets, while retaining memory of all targets referenced in the questions). The underlying reason is that performing complex tasks necessitates the retention of previously explored information across all modules, yet the planner-centric framework fails to facilitate direct interaction between the memory module and other modules. As shown in Figure 1, in the question “Are the sofas in the living room and the media room the same color?”, the agent must explore to multiple objects (“sofa in the living room”, “sofa in the media room”) for comparative reasoning. In the planner-centric framework, due to the lack of interaction between the memory module and planner, the effective information gathered during exploration remain invisible to planner. Specifically, after obtaining the knowledge that “the sofa in the living room is white”, the planner still inefficiently re-explore the same white sofa or fail the task entirely, as this critical input is unavailable during subsequent decision-making.
In this paper, we proposed MemoryEQA, a memory-centric EQA framework (as shown in the right of Figure 2) that maintains a memory database and infuses memory information into all modules, as shown in Figure 1. Unlike the planner-centric methods, where the planner relies only on semantic map, but semantic map is unable to feed into other modules, the proposed memory-centric framework injects memory information into various modules to guide reasoning, enabling efficient exploration and accurate responses.
To achieve this goal, we need to overcome two challenges when designing the memory module. (1) A single form of memory cannot meet the needs of different modules. For instance, the planner requires continuous semantic values (the use of high-dimensional embeddings to model semantic similarity through geometric proximity in a latent space) to guide the next planning objective, while the answering module necessitates observations and historical nature language dialogues pertinent to the query. (2) Natural language memory is difficult to integrate with semantic map and challenging to maintain. Semantic map is structured data (e.g., coordinates, attribute tables), whereas natural language exists as linear sequences, requiring complex conversions (such as mapping “sofa coordinates (x,y)” to “the white sofa on the right”).
To solve the above challenges, we design a multi-modal hierarchical memory, with a corresponding memory storage, retrieval, updating, and injection mechanism. Specifically, multi-modal hierarchical memory comprises global and local memory. Global memory is constructed by combining semantic map with information from object detectors, depth maps, and the agent’s current state, capturing entity locations and the agent’s observation. The local memory stores the current observations, captions of observations obtained through the vision-language model (VLM), and the agent’s current state. For the planner, we project the candidate directions for the next step into pixel space and annotate these candidate directions on the current RGB observation. The memory information is then incorporated as contextual prompts for the VLM to select an optimal direction from the candidates. Regarding the answering and stopping modules, we retrieve the most memory entries relevant to the given question from the memory bank and utilize them as contextual inputs for the VLM. Consequently, natural language serves as an intermediary to adapt memory information to different functional modules. Furthermore, during the construction of natural language-based memory, we employ a detector to localize objects within the scene, followed by leveraging the VLM to extract fine-grained descriptions of these objects. This approach not only enhances the semantic map with natural language annotations but also augments the natural language memory with world coordinate information, thereby enabling integration between natural language memory and the semantic map. Additionally, we have meticulously designed a memory update mechanism to ensure the real-time and relevance of the memory information.
To evaluate memory ability of EQA models, we constructed the MT-HM3D dataset based on HM3D [26], including 1587 QA pairs. Specifically, we utilize the Habitat simulator [30] to sample 5 to 10 images per scene with constrained distances between sampling points. We then prompt GPT-4o [23] to generate questions about multiple targets across different regions, which require the agent to memorize all question-relevant information during exploration. Finally, low-quality data filtering is performed through GPT-4o-mini [23] and manual verification.
Our contributions are as follows:
  • We propose a memory-centric EQA framework that provides a multi-modal hierarchical memory with a corresponding memory retrieval and injection mechanism, and a memory real-time update mechanism for efficient exploration and reasoning.
  • We constructed the MT-HM3D dataset based on the HM3D, offering more complex QA pairs about multiple targets across different regions, effectively evaluating EQA models’ memory capabilities.
  • MemoryEQA achieves an 18.9% improvement in success rate over the baseline model on the MT-HM3D dataset, demonstrating the importance of memory capability in solving complex tasks. Additionally, it sets a new state-of-the-art (SOTA) on the open-vocabulary EQA dataset OpenEQA and achieves comparable success rates to previous best methods on HM-EQA.
Figure 2: Comparison between the planner-centric framework and memory-centric framework.
2 Related Works
2.1 Embodied Question Answering
Embodied question answering [11, 14, 39, 4, 12] has become a challenging task for testing the robot’s ability to autonomously plan tasks and establish semantic understanding of the environment. Das et al. [11] first proposed the EQA task and utilized a planner-controller navigation module to accomplish EQA task in virtual environments. Allen et al. [27] constructed a 2D semantic map of the environment for specific tasks to guide exploration. Answer et al. [2] and Xie et al. [37] developed offline embodied modules that could be queried by large language models (LLMs). Arjun et al. [21] utilized video memory to answer implicit questions using long-context VLMs. Saumya et al. [29] bridged the gap between semantic memory and planners by placing the planner within a compact scene representation. In our work, we position memory as the core module of the embodied question answering framework, enabling the planner, stopping criteria, and question-answering module to reason based on memory information. The memory-centric framework overcomes the limitations of planner-centric methods by infusing memory information into all modules, enabling more efficient exploration and accurate responses through enhanced inter-module interaction.
2.2 Memory Mechanisms in Embodied Agent
Early works utilized the hidden state of the LSTM [15] navigation model as a dense memory representation of the scene [30, 35]. However, the expressive power of this single state vector is limited, making it challenging to handle the storage of memory information in complex 3D real-world scenarios. Recent approaches have modeled scene memory as 2D metric grid [1, 3, 5, 6], topological map [28, 36], and 3D semantic map[9, 24, 31, 29] to serve as global memory information. Although these methods have shown promising performance in navigation task, the complexity of indoor environments, coupled with the lack of local memory information, makes it difficult for models to tackle fine-grained EQA effectively.
Abrar et al. [2] used retrieval-augmented LLM-agent to retrieve relevant memories by forming function calls and answer questions based on a real-time memory-building process. Chen et al. [8] and Xie et al. [37] employed retrieval-augmented generation (RAG) to retrieve relevant images from accumulated observations and used a VLM for inference to generate responses. However, planning that does not rely on memory information will lead to repetitive exploration, resulting in the agent’s inability to answering questions correctly. To this end, we propose a memory-centric EQA framework, MemoryEQA, which maintains a hierarchical memory and flexibly injects memory information into various modules to achieve efficient exploration and accurate responses.
3 Memory-Centric EQA Framework
The overall framework is illustrated in Figure 3, with the memory module in its center, injecting memory information to other modules for reasoning. We define the agent observation sequence as $O=\{{o_{i}}\}_{i=0}^{N_{t}}$, state sequence as $S=\{{s_{i}}\}_{i=0}^{N_{t}}$, and decision sequence as $D=\{{d_{i}}\}_{i=0}^{N_{t}}$, where $N_{t}$ represents the total number of exploration steps. The input question is denoted as $q$. The hierarchical memory module is =$\mathcal{M}=\{M_{j}\}_{j=0}^{N_{s}}$, where $N_{s}$ is the number of scenes. The memory retriever is $\mathcal{F}_{r}$, the answering module is $\mathcal{F}_{a}$, the planner is $\mathcal{F}_{p}$, and the stopping module is $\mathcal{F}_{s}$.
Figure 3: The overall framework of memory-centric EQA. Given the input question and agent observation, we first update the hierarchical memory. Subsequently, question-relevant memories are retrieved through the memory retriever and injected into the respective modules. We concatenate the memories and the question, which are then fed into the planner to determine the next navigation point. Once the agent arrives at a navigation point, it assesses whether to halt its journey. If the decision is ‘No’, the process loops back to the memory update step; if ‘Yes’, the agent responds to the question.
We take the current observations and the question as input, first to update the memory, and then use $\mathcal{F}_{r}$ to retrieve the required memory information. Next, we utilize $\mathcal{F}_{s}$ to determine whether the exploration should be stopped. If sufficient information is available to answer the question, the retrieved memory, the question, and the current state are fed into $\mathcal{F}_{a}$ to generate a response. If the information is insufficient, the retrieved memory, question, state, and observations are input into the planner $\mathcal{F}_{p}$ to derive the next state $s_{i+1}$. In summary, we formulate our framework as
{$r&=\mathcal{F}_{a}(q,s_{t},\mathcal{F}_{r}(q,\mathcal{M})),\quad if\ \ \mathcal{F}_{s}(q,\mathcal{F}_{r}(q,\mathcal{M}))=\text{Yes},\\ s_{i+1}&=\mathcal{F}_{p}(s_{i},o_{i},q,\mathcal{F}_{r}(q,\mathcal{M})),\quad else,\end{aligned}\right.$ (1)
where $r$ represents the agent response and $s_{t}$ is the status when the agent stops exploring.
Figure 4: Hierarchical Memory. The global memory (yellow box) is constructed by annotating entity states on a real-time generated map, and local memory consist of he visual observation and object information (gray box). All memories are persisted and used as a database for scene retrieval (green box).
3.1 Planner
We employ a dynamic 3D voxel-based mapping approach for robotic exploration. The environment is represented using an extensible voxel grid with initial dimensions $L\times W\times H$, where height $H$ remains fixed at $3.5m$ (typical floor height) while length $L$ and width $W$ dynamically expand as exploration progresses. Each cubic voxel with edge length $l$ is updated in real-time through Truncated Signed Distance Function (TSDF) fusion[22, 40]: upon acquiring depth image, the system first updates occupancy states for all visible voxels while selectively marking exploration status only for voxels within a constrained field of view.
For path planning purposes, the 3D voxel map is projected onto a 2D point map M, where each point’s status is determined by two criteria: (1) it is marked as traversable when all voxels below camera elevation height are free space, and (2) considered explored only when all voxels along the vertical dimension have been explored. Building upon this representation, the system implements frontier-based exploration [38] for autonomous navigation by detecting boundaries between explored and unexplored regions, sampling candidate poses from these frontiers, and determining optimal exploration directions to achieve efficient environment mapping.
3.2 Hierarchical Memory
As shown in Figure 4, our hierarchical memory is divided into local memory and global memory. The corresponding memory storage, retriever, injector, and memory update mechanism collectively form a closed loop of “construction-utilization-updating”.
Local Memory.
Local memory is composed of visual observations, goal descriptions, planning information, and the agent’s current state. Specifically, based on the current visual observation, we use an object detector to identify the objects present in the current view, and obtain fine-grained descriptions of these objects by prompting a VLM. Planning information includes the current planning step and the decision (derived from the offset between the navigation point and the current position). The agent’s current state encompasses its location, orientation, and a coarse-grained description of the space it is in (obtained by prompting the VLM for a description of the current observation).
Gloabl Memory.
Building upon the real-time constructed map, the global memory annotates room information and target information. Among these, room information primarily includes the room category and its location (represented by the position from which the agent observes the room). Target information encompasses the target’s position, category, description, as well as the position and orientation from which the agent observes the target. Details on prompts about memory construction can be found in the supplemental materials.
3.2.1 Memory Storage
The memory information is stored in dense vectors to facilitate efficient retrieval. We employ a unified encoder $e_{c}(\cdot)$ to extract multi-modal memory features $f_{i}=e_{c}(M_{i})$. These feature vectors are then stored in a vector library =$\mathcal{L}=\{L_{i}\}_{k=0}^{N_{s}}=\{<I_{i},M_{i},f_{i}>\}_{i=0}^{N_{s}}$, with each memory assigned a unique index $I_{i}$ for subsequent retrieval operations.
3.2.2 Memory Retriever
Before the reasoning process within each module, it is essential to retrieve the necessary memories from the memory database. To this end, we have designed a specialized retriever tailored for the hierarchical memory retrieval system. Our memory retriever $\mathcal{F}_{r}$ consists of two components:
Scene Retrieval.
Scene retrieval occurs only when the agent initializes. Upon obtaining the initial observation $o_{0}$, we extract its feature $f_{o_{0}}=e_{c}(o_{0})$. We compute cosine similarity between $f_{o_{0}}$ and vectors $\mathcal{L}$ in the vector library, and select the $\text{top}_{k}$ most similar features (requiring similarity scores to exceed a threshold)
$I^{s}=\text{Index}_{\mathcal{L}}(\text{top}_{k}(\text{sim}(f_{o_{0}},\mathcal{L}))),$ (2)
where $I^{s}$ is the set of scene index, and $\text{Index}_{\mathcal{L}}(\cdot)$ is the operator of index in $\mathcal{M}$. Next, we identify the scene index $I^{c}=\text{Max}(\text{Fre}(I^{s}))$ with the highest number of occurrences as the current scene, $\text{Fre}(\cdot)$ is the operator used to calculate the frequency of occurrence of elements. The vectors associated with this scene form a sub-database for subsequent retrieval steps.
Content Retrieval.
For content retrieval, we adopt a multi-modal retrieval approach. Given the current observation $o_{i}$ and user question $Q$, we use a unified encoder to generate a combined feature
$f_{q}=\text{Norm}(\operatorname{concat}(e_{c}(o_{i}),e_{c}(q))).$ (3)
We then compute similarity between $f_{q}$ and vectors in the scene memory $M_{i}$ (from the result of scene retrieval and prior exploration) and retrieve the $\text{top}_{k}$ most similar indices as the required memory content. We final get the local memory
$\begin{split}\mathcal{M}^{s}=\text{top}_{k}(\{M_{i}|\|(f_{q}-M_{i}\|_{2}<\alpha_{e}(1+\text{Entropy}(f_{q})))\cap\text{sim}(f_{q},M_{i})>\alpha_{s}\}),\end{split}$ (4)
where $\mathcal{M}^{s}$ represents the memory required in the current state, $\alpha_{e}$ and $\alpha_{s}$ are the weight of similarity. We utilize entropy scaling to adapts the search scope to query complexity and semantic similarity greater than a threshold $\alpha_{s}$.
The size of memory significantly impacts model performance (see supplementary material). To address this, we employ a dynamic k-value mechanism that adaptively samples varying amounts of memory information according to different scenarios. Specifically, we implement an entropy-based linear adjustment strategy formulated as:
$k=\left\lceil k_{\min}+\beta\cdot\text{Entropy}(f_{q})\right\rceil,$ (5)
[…]
3.2.3 Memory Injector
The injection of global memory only needs to be loaded before the agent makes its first step of exploration. For local memory, for the planner, we map $n$ candidate exploration directions to pixel space, marking the candidate directions in the current observed RGB image to obtain $d_{i}$. We use $d_{i}$, memory, and the question as inputs and prompt the VLM to select one of the candidate directions. Based on the VLM’s choice, we construct a one-hot vector as the semantic weight for the planner. For the answering module, we directly input the current observation $o_{i}$, memory, and the question as prompts to the VLM to obtain the response.
3.2.4 Memory Update
After loading the global memory, continued exploration may reveal new environmental changes that necessitate memory updates. To address this, we have designed a set of memory update mechanism to effectively prevent erroneous triggering of updates. The specific rules are as follows:
  • Agent enters novel region with
    $\min_{m_{j}\in M}\|P_{i}-P_{j}\|_{2}>\beta_{p},$ (6)
    $\min_{m_{j}\in M}\angle(R_{i},R_{j})>\beta_{r},$ (7)
    where $M$ is the set of memories. $P_{i}$ and $R_{i}$ is the $i$-th position and rotation in memories. $\|\cdot\|_{2}$ is the Euclidean distance operator. (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ ) is the Euclidean angle operator. In other words, the distance between the current position and orientation and all the positions in memory is greater than threshold $\beta_{p}$ and $\beta_{r}$.
  • The similarity between the current observation and the memory observation is below a threshold. We calculate similarity
    $SIM(o_{i},o_{j})=\alpha\cdot\text{SSIM}(o_{i},o_{j})+(1-\alpha)\cdot\text{sim}(f_{i},f_{j}),$ (8)
    by integrating both structural similarity and semantic similarity, where $(f_{i},f_{j})=(e_{c}(o_{i}),e_{c}(o_{j}))$ is current observation feature and history observation feature respectively. The SSIM [33] is structural similarity.
  • The current observation’s field of view is not restricted (black areas do not exceed half of the observation).
The memory update mechanism is triggered only when all the above rules are satisfied. Due to the nature of our memory storage, updating memory can be efficiently achieved by modifying the corresponding memory index.
3.3 Discrete Stop Criterion
We adopt a discrete stopping criterion, where we set a multiple-choice question and let the model assess its confidence level in answering the question. If the confidence exceeds a given threshold, the exploration stops, and the answering module is invoked. If the maximum number of steps is reached, the agent will also stop exploring and provide an answer. So the probability modeling of the stop criterion is $\mathcal{F}_{s}=\mathbf{1}(\mathcal{P}(S|M_{i})>\gamma),$ where $\mathbf{1}$ is an indicator function, which is 1 when the condition is met and 0 otherwise. $\gamma$ is threshold, and $S$ means current stats of the agent. In the implementation, we prompt VLMs to provide the confidence level in answering the question based on the retrieved memory information from the current observation. For detailed information, please refer to the supplementary material.
3.4 Answering Module
We employ a retrieval-augmented VLM as the answering module, which is invoked after the agent terminates exploration. By selectively feeding only memory content relevant to the query into the VLM, we achieve both computational efficiency and reduced interference from extraneous information.
4 Data Construction
To evaluate memory ability of the models, we propose an multi-target EQA dataset based on HM3D [26]. Our dataset includes 1587 QA samples on 500 scenes (more statistical data see the supplemental material). The most significant distinction from existing datasets lies in the fact that the MT-HM3D dataset constructs question-answering pairs targeting multiple regions and multiple objects in real-world scenarios, while exhibiting a richer diversity of question formats. We define four types of questions:
  • Comparison: comparing status two or more entities, e.g., “Is the bed closer to the chair than to the window in the bedroom?”
  • Relationship: asking about the relationship between two entities, e.g., “Are the red lampshade and the pipes in the same room?”
  • Counting: requiring the agent to count an entity, e.g., “How many chairs are around the dining table?”
  • Attribute: assessing the difference in attributes between similar entities, e.g., “Is the table in the kitchen bigger than the table in the study?”
We use GPT-4o [23] and GPT-4-mini [23] to construct multi-target question answering. The process consists of the following steps: (1) Sample 5 to 10 images from the scene using the Habitat simulator. (2) Input the sampled images and prompt the GPT-4o [23] select multiple related objects from those images, specifying the output format to obtain raw data. (3) Use GPT-4-mini [23] to filter the raw data, removing illogical question-answer pairs, and generate diverse QA data. (4) Manually verify the correctness of the data to ensure that the questions can be answered within the embodied environment. Specific prompts and more details are provided in the supplementary material.
Metrics. We evaluate the embodied question-answering model using success rate (%), normalized steps, GPT score, and ROUGEL [19] (the details see supplemental material).
5 Experiments
5.1 Implementation
We use QwenVL2-7B [32] and GPT-4o [23] as the built-in multi-modal large models, and employ Clip-ViT-Large [25] as unified encoder to encode the multi-modal memory into 768-dimensional features. The object detector is YOLOv11[18]. Our simulator follows the setup of Explore-EQA [27] and conducts experiments in Habitat-Sim [30], and we conduct real-world experiments using the Unitree Go2 robot (see supplementary material). Inference is performed on an NVIDIA Tesla L40, with approximately 40 hours required to complete inference on the MT-HM3D dataset. We use Faiss [13] as a library of dense vectors.
Table 1: Experiments on multiple methods across MT-HM3D, HM-EQA, and OpenEQA, testing various foundational models. denotes that the value was converted based on the metrics provided in the original paper. $\ddagger$ indicates that the performance was obtained through our replication.
Exp. Model VLM MT-HM3D HM-EQA OpenEQA
Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$ GPT $\uparrow$ ROUGEL $\uparrow$ Step $\downarrow$
1 ExploreEQA Prismatic-7B 35.1 0.64 51.5 0.75 28.28 0.26 0.695
2 Qwen2VL-7B 36.1 0.66 44.31 0.74 27.36 0.241 0.741
3 GPT4o 36.2 0.64 47.4 0.77 29.21 0.255 0.773
4 GraphEQA Qwen2VL-7B 45.63 0.45 56.7 0.35 30.1 0.277 0.625
5 GPT4o - - 63.5 0.2 - - -
6 MemoryEQA Qwen2VL-7B 51.2 0.4 59.9 0.52 34.64 0.3 0.6
7 GPT4o 55.1 0.41 63.4 0.4 36.8 0.32 0.552
5.2 Main Results
As shown in Table 1, we evaluate three approaches ExploreEQA [27], GraphEQA [29], and MemoryEQA across the MT-HM3D, HM-EQA [27], and OpenEQA [21] datasets, testing their compatibility with diverse VLMs (Prismatic [17], Qwen2VL [32] and GPT4o [23]).
On MT-HM3D, MemoryEQA attains a success rate of 54.5%, outperforming baseline by 18.9% (Exp.3), highlighting the critical role of hierarchical memory in multi-target tasks. The results demonstrate that MemoryEQA exhibits superior performance in multi-modal reasoning tasks, particularly in complex scene understanding and knowledge integration.
The analysis of experimental results across Exp.1-3, Exp.4-5, and Exp.6-7 reveals a significant positive correlation between the performance of the VLM and the effectiveness of the EQA system. This observation underscores the critical role that VLM plays in enhancing the EQA system’s ability to process and interpret complex queries within visual environments. As the VLM’s accuracy and understanding improve, so does the EQA system’s capacity to deliver precise and contextually relevant answers, demonstrating a synergistic relationship between the two components. This finding highlights the importance of advancing VLM capabilities to further boost the overall performance of EQA systems in practical applications.
Table 2: Ablation study on MT-HM3D and HM-EQA to evaluate the impact of injecting memory into different modules on performance. (S: Stop Criterion, A: Answering Module, P: Planner)
Module MT-HM3D HM-EQA
Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$
None 34.3 0.7 44.6 0.79
S 35.1 0.66 43.5 0.7
S + A 46.8 0.69 52.2 0.71
S + A + P 51.2 0.4 59.9 0.52
Table 3: The impact of global memory and local memory on MemoryEQA.
Method Succ. $\uparrow$ Step $\downarrow$
ExploreEQA 51.5 0.75
MemoryEQA w/o Global Memory 56.4 0.74
MemoryEQA w/o Local Memory 52.15 0.6
MemoryEQA 59.9 0.52
5.3 Ablation Studies
As shown in Table 2, the ablation study evaluates the impact of the injection of memory into various module on MT-HM3D and HM-EQA datasets. Without memory injection (None), baseline success rates are 34.3% (MT-HM3D) and 44.6% (HM-EQA), with suboptimal step efficiency. Introducing the Stop Criterion (S) alone yields marginal improvements in MT-HM3D (+0.8% Succ., reduced steps). Combining Stopping and Answering Module (S+A) significantly boosts performance: MT-HM3D’s success rises by 12.5% (to 46.8%), and HM-EQA improves by 7.6% (to 52.2%), confirming that memory-enhanced answering critically resolves ambiguities in both tasks. Full integration (S+A+P) achieves peak results: MT-HM3D reaches 51.2% Succ. (16.9% gain over baseline) with a 0.43 norm step reduction, while HM-EQA attains 59.9% success rate (+15.3%) and 25% fewer steps (0.52). The planner’s contribution is crucial, in synergy with S and A to optimize task decomposition and trajectory planning. This demonstrates that memory injection benefits all modules cumulatively, with the planner driving efficiency by mitigating redundant exploration. Notably, HM-EQA’s higher baseline and performance with S+A+P suggest its questions demand multi-phase reasoning, where memory-centric framework excels. In summary, the injection of memory into various modules is essential to improving success rates and efficiency in EQA tasks.
We have evaluated the impact of both global and local memory on the model’s performance. As shown in Table 3, both local and global memory play significant roles in problem-solving. Specifically, local memory enhances the success rate by 4.9%, while global memory boosts exploration efficiency by approximately 20%. From these findings, we can conclude that global memory retains the location and category of entities within the scene, aiding the planner in navigating directly to the target location and enabling the answering module to respond to questions based on past explorations. Local memory, on the other hand, preserves fine-grained information during the current exploration, allowing the model to answer questions with greater accuracy.
6 Limitations
In our implementation, the exploration process lacks interpretability and a basis for decision-making. Additionally, the model may generate responses without relying on the information obtained during exploration, and current evaluation metrics may inaccurately assess performance in such cases. In the future, we will build trajectory annotations based on the MT-HM3D, including thought processes, tool-chains, and the outputs of executing these tool-chains. Based on data with trajectory annotations, we will design an evaluation metric that takes the exploration process into account.
7 Conclusion
In this paper, we propose a memory-centric EQA framework, which contrasts with planner-centric frameworks. In our method, hierarchical memory information is feed into the planner, stop module, and answering module. The memory guides reasoning for various module, prevented redundant exploration and improved success rate of response in multi-target, multi-scenario tasks. Additionally, we have constructed a multi-target EQA dataset MT-HM3D based on the real-world HM3D scenes. This dataset includes four categories of multi-target tasks, allowing for comprehensive validation of the memory capabilities of EQA models. The effectiveness of MemoryEQA has been demonstrated through performance on the HM-EQA and MT-HM3D datasets.
Supplementary Material
Appendix A Method Details
We utilize prompt 1 and prompt 2 to prompt VLM to generate coarse-grained scene description and fine-grained object description respectively.
Describe this image. Output in the following format:
Room: <room>
Object: <obj1>, <obj2>,...
Description: <description>
Prompt 1: Prompt for coarse-grained scene description.
Please describe the category, arrtibute, and description of the object.
Output in the following format:
cate: [category]
attr: [arrtibute]
desc: [description]
Prompt 2: Prompt for fine-grained object description.
For the stopping module, we prompt the VLM to evaluate the confidence of the response through a multiple-choice question format. If the model selects a confidence level of C or higher, it will cease exploration and proceed to provide the answer.
Consider the question: ‘{Question}’. How confident are you in answering this question from your current perspective?
A. Very low
B. Low
C. Medium
D. High
E. Very high
Answer with the option’s letter from the given choices directly.
Prompt 3: Prompt for the evaluation of confidence.
Prompt 4 and prompt 5 are used to prompt the answering module to closed-vocabulary response and open-vocabulary response.
{Question} Answer with the option’s letter from the given choices directly.’
Prompt 4: Prompt for the response.
{Question} Answer with the brief sentence.
Prompt 5: Prompt for the response on open-vocabulary data.
Next, prompt 6 and prompt 7 follow ExploreEQA [27] setting to prompt VLM to output explore direction.
Consider the question: ’{Question}’, and you will explore the environment for answering it.
Which direction (black letters on the image) would you explore then? Provide reasons and answer with a single letter.
Prompt 6: Prompt for the explore direction.
Consider the question: ’{Question}’, and you will explore the environment for answering it.
Is there any direction shown in the image worth exploring? Answer with Yes or No.
Prompt 7: Prompt for the Exploring Value.
Appendix B Data Construction
Figure 5: Statistic of our proposed MT-HM3D.
B.1 Statistics
To evaluate the memory capabilities of EQA models, we propose an EQA dataset that features multiple targets distributed across various regions. As shown in Table  4, our dataset exhibits the highest average number of targets and regions, significantly increasing the complexity of the task. The requirement to explore multiple targets across different regions poses a substantial challenge to the model’s memory capacity. The statistical details of the MT-HM3D dataset are illustrated in Figure  5.
Table 4: The comparison between our proposed MT-HM3D and existed dataset.
Dataset Simulator Source Real Scenes Objects Region Creation
EQA-v1[11] House3D SUNCG - - Rule-Based
MP3D-EQA[34] MINOS MP3D - - Rule-Based
MT-EQA[39] House3D SUNCG - - Rule-Based
IQA[14] AI2THOR - - - Rule-Based
EXPRESS-Bench[16] Habitat HM3D - - VLMs
HM-EQA[27] Habitat HM3D 2.87 0.51 VLMs
Open-EQA[21] Habitat ScanNet/HM3D 2.53 0.27 Manual
MT-HM3D (Ours) Habitat HM3D 3.66 0.947 VLMs
B.2 Metrics
For multiple-choice questions, we have the model directly provide a predicted option, and then calculate the success rate (%). For open-vocabulary questions, we use two metrics to evaluate performance of the model: GPT score and ROUGEL [19]. The specific calculation methods are as follows:
$ROUGE_{L}=\frac{2LCS(C,R)}{|R|+|C|},$ (9)
$S_{GPT}=GPT4o(promtp,C,R),$ (10)
where $LCS(\cdot,\cdot)$ is longest common subsequence, $C$ is prediction sentence, $R$ is ground-truth sentence, $cos$ is cosine similarity, ||$\cdot$|| ⋅ | is calculating the number of words.
And then we use average normalization steps to evaluate the efficiency of the agents. The specific calculation formula is
$Norm_{step}=\frac{1}{N}\sum_{i=0}^{N}\frac{N_{i}}{\sqrt{S_{i}}*\gamma_{s}},$ (11)
where $N$ means the number of samples, $N_{i}$ represent the explore step number of agent on current sample, $S_{i}$ is room size, and $\gamma_{s}$ is the relation ratio between max steps and room size.
B.3 Prompts
We sample multiple images from the scene, stitch them together, and input the images along with a prompt 8 to generate proposed question-answer pairs.
Select one or some object from those image (they must be related objects).
Now, I hope you POSE A QUESTION based on the relationship between these two objects, and the question must meet the following requirements:
1. Select the question label from the following list: [Comparison, Relationship, Counting, Attribute];
1. do not ask "why" and "how";
2. the names of the objects must appear in the question;
3. avoid descriptions such as "left image" or "right image";
4. provide four candidate answer options as well as the correct answer option.
For example:
Does <OBJ1> share same color as <OBJ2> in <ROOM>?
Does <OBJ1> in <ROOM1> share same color as <OBJ2> in <ROOM2>?
Is <OBJ1> bigger/smaller than <OBJ2> in <ROOM>?
Is <OBJ1> in <ROOM1> bigger/smaller than <OBJ2> in <ROOM2>?
Is <OBJ1> closer than/farther from <OBJ2> than <OBJ3> in <ROOM>?
Is <ROOM1> bigger/smaller than <ROOM2> in the house?
How many <OBJ1> in <ROOM1>?
How many rooms have <OBJ1>?
The output must strictly following the format below:
Objects: [object1]; [object2]
Question: [question]
Options: [A. xxx; B. xxx; C. xxx; D. xxx]
Answer: [answer]
Label: [label]
Prompt 8: Prompt for data generation.
After obtaining the question-answer pair, we use prompts 9 to verify the validity of the question and answer.
I need you to help me check whether the question meets my requirements, which are as follows:
(1) The question requires additional visual input to answer;
(2) The objects mentioned in the question can be found in an indoor scene;
(3) For an exploration robot in any location, the references in the question are clear during the robot’s autonomous exploration process.
The question is as follows: {}
Please help me check whether the question meets the requirements. If it does not meet the requirements, answer "No" and provide a reason; if it meets the requirements, simply answer "Yes."’
Prompt 9: Prompt for Verify the rationality of the data.
Appendix C More Experiments
Table 5: All hyper-parameters list.
Hyperparameter Value Hyperparameter Value
seed 42 device cuda
vlm
model name or path Qwen2-VL-7B
rag
use rag true text clip-vit-large-patch14
visual clip-vit-large-patch14 dim 1536
max retrieval num 10
camera/image
detector yolo11x.pt camera height 1.5
camera tilt deg -30 img width 640
img height 480 hfov 120
tsdf grid size 0.1 margin w ratio 0.25
margin h ratio 0.6
navigation
init clearance 0.5 max step room size ratio 3
black pixel ratio 0.7 min random init steps 2
planner
dist T 10 unexplored T 0.2
unoccupied T 2.0 val T 0.5
val dir T 0.5 max val check 3
smooth sigma 5 eps 1
min dist from cur 0.5 max dist from cur 3
frontier spacing 1.5 min neighbors 3
max neighbors 4 max unexplored 3
max unoccupied 1
visual prompt
cluster threshold 1.0 num prompt points 3
num max unoccupied 300 min points clustering 3
point min dist 2 point max dist 10
cam offset 0.6 min prompt points 2
circle radius 18
C.1 Implementation Details
We have listed all the hyper-parameters used for our best results in Table  5.
C.2 Parameter Analysis
As shown in Table  6, the selection of the K value is a pivotal factor, as both excessively large and unduly small K values can precipitate suboptimal performance. An insufficiently small K value engenders a dearth of memorized information, whereas an inordinately large K value not only induces informational redundancy but also fosters hallucinations in the response module. We believe that using different K values for different problems may be a more optimal strategy. Specifically, a dynamic memory window, rather than a fixed K value, could better adapt to varying complexities, which warrants further investigation.
Table 6: The impact of memory retrieval quantity on model performance.
K Succ. $\uparrow$ Step $\downarrow$
0 34.3 0.7
2 44.3 0.45
4 49.65 0.4
8 48.7 0.42
12 46.21 0.3
C.3 Real-World Results
We conducted experiments using the Unitree Go2 robot in both office settings and artificially constructed scenarios. The RGB-D images were captured using the RealSense D435i depth camera, with DDS serving as the underlying middleware, and the control of the agent was accomplished on the ROS2 platform [20]. The multi-modal large model was deployed on 8 * Nvidia Tesla L40.
As shown in Figure  6, in the handmade scenario, we posed an open-vocabulary question to the agent: “How many apples are there in the scene?" After a thorough exploration, the agent provided an answer and did not recount when revisiting previously explored areas. In the office setting, we inquired, “Is the black garment placed on the chair?" The agent swiftly deduced the correct answer following a simple planning process.
Figure 6: Experiment on real world. Left: handmade scenario. Right: office
C.4 Visualization
Figure 7: Visualization of Two-Round QA in the HM3D Dataset. The figure above illustrates the performance of the first round of QA baseline and MemoryEQA, with red dots representing the planning points of ExplorEQA and green lines indicating the planning path of MemoryEQA. The figure below demonstrates the results of the second round of QA, where the information collected during the exploration process of the first round is used as memory for further planning.
As shown in Figure  7, we have visualized ExploreEQA and MemoryEQA to demonstrate the advantages of our global and local memory mechanisms. The figure above illustrates an example of local memory. The baseline method provided an incorrect answer immediately after observing a sofa, whereas MemoryEQA, upon observing the first sofa, did not retrieve information about another sofa from its memory and thus continued its exploration, ultimately delivering the correct response.
We stored information collected during the exploration process depicted in the figure above as persistent memory and input it into the question-answering process shown in the figure below. Since the yellow sofa was observed at the 10-th step during the first question, the agent, upon revisiting the scene, could prioritize exploration based on the coordinates of the yellow sofa observed previously, requiring only four steps to provide the correct response. In contrast, the baseline method took 14 steps of exploration to arrive at the correct answer.